# Kazhdan groups of dimension $16$ with an uncountable range of second $\ell^{2}$-Betti numbers

### v1

@@ 4.3 Lacunary hyperbolic groups
Let $G_{1}\to G_{2}\to\cdots$ be a directed sequence of epimorphisms of hyperbolic groups, with colimit $G_{\infty}$. Let $S_{1}$ be a finite generating set for $G_{1}$, and let $S_{i}$ be its pushforward to $G_{i}$; let $\delta_{i}$ be the hyperbolicity constant of $\mathrm{Cay}(G_{i},S_{i})$. For each $i\in\mathbb{N}$, let $\rho_{i}$ be the radius of the largest ball centred at the identity on which the quotient $G_{i}\to G_{i+1}$ is injective.
@@
5 Failure of semi-continuity
In this section we prove C. We start with a more precise statement.
@@ 5 Failure of semi-continuity
Let $G$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, and let $B\subset G$ be a finite subset. Then for every $n\in\mathbb{N}$, there exists a torsion-free hyperbolic quotient $G\to G_{n}$, injective on $B$, of cohomological dimension $16$, such that for every unitary $G_{n}$-representation $V$, the second homology is $H_{2}(G_{n};V)\cong V^{n}$.
Cocompact lattices in the isometry group of the octonionic hyperbolic plane exist [BHC62], and they can be chosen torsion-free up to passing to a finite index subgroup [Sel60].
@@
The group $G$ is torsion-free hyperbolic of cohomological dimension $16$, and by Theorem 3.10 it has property $T_{2}]$. We construct $G_{n}$ by induction on $n$, starting with $G_{0}=G$, which works by Proposition 3.8.
Suppose by induction that $G_{n-1}$ has been constructed. Lemma 4.9 gives an element $g\in G_{n-1}$ and we define $G_{n}$ to be the corresponding group-theoretic Dehn filling. Lemma 4.9 gives injectivity on $B$ and the fact that $G_{n}$ is torsion-free hyperbolic. The first part of Corollary 4.3 implies that the cohomological dimension of $G_{n}$ is $16$. Moreover, since $G_{n-1}$ has property $(T)$, applying Proposition 3.8 and the second part of Corollary 4.3, we obtain, for every unitary $G_{n}$-representation $V$, a short exact sequence
$0\to H_{2}(G_{n-1};V)\to H_{2}(G_{n};V)\to V\to 0.$
By the induction hypothesis the first term is $V^{n-1}$. Therefore $H_{2}(G_{n};V)\cong V^{n}$, which proves the claim and hence the theorem. ∎
@@
The groups $G_{n}$ of course fail to have property $(T_{2})$, despite being quotients of the $(T_{2})$ group $G$, as remarked in [FF25]. Still, the second homology (and similarly the second cohomology) functor enjoys some strong form of rigidity: in a way, its output can be read directly from its input. It would be interesting to explore whether a (co)homological rigidity of this form characterises quotients of $(T_{2})$ groups.
@@
We have $b^{(2)}_{2}(G)=0$ by Corollary 3.9. For each $i\in\mathbb{N}$, we let $B_{i}$ be the ball of radius $i$ in $G$, for a fixed choice of finite generating set, and $n_{i}\in\mathbb{N}$ as in the statement of C. Then Theorem 5.1 builds a group $G_{i}$, which by Proposition 3.6 has
$b^{(2)}_{2}(G_{i})=\dim_{\mathcal{N}(G_{i})}H_{2}(G_{i};\ell^{2}(G_{i}))=\dim_{\mathcal{N}(G_{i})}\ell^{2}(G_{i})^{n_{i}}=n_{i}.$
@@
In Theorem 5.1, we used the full power of property $T_{2}]$, but the proof shows that if $V$ is assumed to have no non-zero invariant vectors, then property $(T_{2})$ suffices. In the proof of C, we applied Theorem 5.1 only with $V=\ell^{2}(G_{i})$, which has no non-zero invariant vectors, so property $(T_{2})$ would have been sufficient. Similarly, the computation of second $\ell^{2}$-Betti numbers in the next section could be obtained only using property $(T_{2})$.
6 Proof of the main theorem
@@
Let $p$ be a prime. There exists an uncountable family of groups $(G_{\alpha})_{\alpha\in A}$ and an integer $\chi\geq 1$ with the following properties.
@@
    $G_{\alpha}$ has property $(T)$ and is lacunary hyperbolic.
@@
    $G_{\alpha}$ has cohomological dimension $16$ modulo $p$.
@@
    The second $\ell^{2}$-Betti numbers $b^{(2)}_{2}(G_{\alpha})$ are pairwise distinct.
@@
    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{\alpha})=b_{2}^{(2)}(G_{\alpha})+\chi$.
@@
We choose as an index set $A\coloneqq\{2,3\}^{\mathbb{N}}$, and fix an element $\alpha\in A$. The group $G_{\alpha}$ will arise as the colimit of a direct sequence of epimorphisms of hyperbolic groups $G_{1}\to G_{2}\to\cdots$, where $G_{1}$ is a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, which has Euler characteristic $\chi\geqslant 1$ [Lüc02, Corollary 5.16(1) on p. 231]. We fix a finite generating set $S_{1}$ for $G_{1}$ and let $S_{i}$ be the pushforward to $G_{i}$. We also specify elements $g_{i}\in G_{i}$, and will prove the following additional properties.
@@
    The element $g_{i}\in G_{i}$ has infinite order, and for all $j>i$ its image in $G_{j}$ has order $p^{a_{i}}$, where $a_{i}\in\mathbb{N}$ is a sequence with the property that $a_{i+1}-a_{i}-1$ is a positive power of $\alpha(i)\in\{2,3\}$.
@@
Let us first prove that the colimit $G_{\alpha}$ has the desired properties. The group $G_{1}$ has property $(T)$ by Theorem 3.10, hence so does $G_{\alpha}$, being its quotient. Item 2 implies that $G_{\alpha}$ is lacunary hyperbolic, by Theorem 4.10. Now $G_{1}$ is the fundamental group of a closed $16$-dimensional aspherical manifold, so it has cohomological dimension $16$, and $H^{16}(G_{1};R)\cong R$ for every ring $R$. Therefore Item 3 and Corollary 2.4 imply that $G_{\alpha}$ has cohomological dimension $16$ modulo $p$.
Next, we compute $b^{(2)}_{2}(G_{\alpha})$. By Proposition 2.2, we have
$H_{2}(G_{\alpha};\ell^{2}(G_{\alpha}))\cong\bigoplus_{i\geqslant 1}\ell^{2}(G_{\alpha})_{g_{i}}.$
Each $g_{i}$ has order $p^{a_{i}}$ in $G_{\alpha}$, hence by Proposition 3.6 and Proposition 3.11:
$\dim_{\mathcal{N}(G_{\alpha})}\ell^{2}(G_{\alpha})_{g_{i}}$ $=\dim_{\mathcal{N}(G_{\alpha})}H_{0}(\langle g_{i}\rangle,\ell^{2}(G_{\alpha}))=\dim_{\mathcal{N}(G_{\alpha})}H_{0}(\langle g_{i}\rangle,\mathcal{N}(G_{\alpha}))$
$=\dim_{\mathcal{N}(\langle g_{i}\rangle)}H_{0}(\langle g_{i}\rangle,\mathcal{N}(\langle g_{i}\rangle))=b^{(2)}_{0}(\langle g_{i}\rangle)=p^{-a_{i}}.$
@@
$b^{(2)}_{2}(G_{\alpha})=\dim_{\mathcal{N}(G_{\alpha})}H_{2}(G_{\alpha};\ell^{2}(G_{\alpha}))=\sum\limits_{i\geqslant 1}\frac{1}{p^{a_{i}}}.$
Reading this number in base $p$, we see a $1$ at position $a_{i}$, then a string of 0s of length $(a_{i+1}-a_{i}-1)$ before the next $1$ at position $a_{i+1}$. The length of this string of 0s is either a power of $2$ or a power of $3$ depending on whether $\alpha(i)=2$ or $3$. This shows that we can read $\alpha$ from $b^{(2)}_{2}(G_{\alpha})$.
Finally, we compute $\chi^{(2)}(G_{\alpha})$. By Item 3 and Proposition 2.2, for all $n\geqslant 3$ we have
$b^{(2)}_{n}(G_{\alpha})=\dim_{\mathcal{N}(G_{\alpha})}H_{n}(G_{\alpha};\mathcal{N}(G_{\alpha}))=\dim_{\mathcal{N}(G_{\alpha})}H_{n}(G_{1};\mathcal{N}(G_{\alpha}))=b^{(2)}_{n}(G_{1}\to G_{\alpha}).$
Since $G_{\alpha}$ has property $(T)$, $b^{(2)}_{1}(G_{\alpha})=0$ by [BV97, Corollary 6]. Hence
$\chi^{(2)}(G_{\alpha})$ $=\sum\limits_{n\in\mathbb{N}}(-1)^{n}b^{(2)}_{n}(G_{\alpha})=b^{(2)}_{2}(G_{\alpha})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{\alpha})$
$=b^{(2)}_{2}(G_{\alpha})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{1}\to G_{\alpha})=b^{(2)}_{2}(G_{\alpha})+\chi,$
@@
[…]
By Proposition 4.7 again, the triple $(G_{i},\langle g_{i}\rangle,\langle g_{i}^{k}\rangle)$ is Cohen–Lyndon. Because $k$ is a power of $p$, for every ring $R$ in which $p$ is invertible, we can apply Corollary 4.3. The first part immediately gives Item 3. Moreover, the second part, together with the fact that $G_{i}$ has property $(T)$, gives a short exact sequence
@@
We now prove A. The proof will be similar, but more technical, in that we need to leverage an additional prime, and the full power of Proposition 4.8, to make the group $G_{\alpha}$ simple. This improvement employs a trick from [OT13], where it was used to construct simple groups with large first $\ell^{2}$-Betti numbers. The following will be our starting point.
@@
Let $G$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane. Let $q$ be a prime. Then there exists a quotient $G\to G_{1}$ with the following properties.
@@
    $G_{1}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
@@
    The quotient $G\to G_{1}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{1}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible.
@@
    $G_{1}$ is generated by elements $\{x_{1},\ldots,x_{m}\}$, each of which has order a power of $q$.
@@
    There is a finite set $\mathcal{E}$ of finite subgroups of $G_{1}$, such that for every unitary $G_{1}$-representation $V$, the second homology is
    $H_{2}(G_{1};V)=\bigoplus_{E\in\mathcal{E}}V_{E}.$
@@
We start by choosing a generating set $\{x_{1}^{0},\ldots,x_{m}^{0}\}$ of $G$. For $i=0,\ldots,m$, we construct by induction a quotient $G\to G_{i/m}$ with the following properties.
@@
    $G_{i/m}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
@@
    The quotient $G\to G_{i/m}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{i/m}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible. In particular $G_{i/m}$ has cohomological dimension $16$ modulo $q$, and so it only contains $q$-torsion [Bie81, Proposition 4.11 on p. 63].
@@
    $G_{i/m}$ is generated by elements $\{x_{1}^{i},\ldots,x_{m}^{i}\}$, where $x_{1}^{i},\ldots,x_{i}^{i}$ have order a power of $q$.

### v2

@@ 4.3 Lacunary hyperbolic groups
Let $G_{0}\to G_{1}\to\cdots$ be a directed sequence of epimorphisms of hyperbolic groups, with colimit $G_{\infty}$. Let $S_{0}$ be a finite generating set for $G_{0}$, and let $S_{i}$ be its pushforward to $G_{i}$; let $\delta_{i}$ be the hyperbolicity constant of $\mathrm{Cay}(G_{i},S_{i})$. For each $i\in\mathbb{N}$, let $\rho_{i}$ be the radius of the largest ball centred at the identity on which the quotient $G_{i}\to G_{i+1}$ is injective.
@@
5 Rationals and hyperbolic groups
In this section we prove B and C.
5.1 Realising natural numbers
We start with the torsion-free case, which is simpler, and sufficient for C.
@@ 5.1 Realising natural numbers
Let $G$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, and let $B\subset G$ be a finite subset. Then for every $m\in\mathbb{N}$, there exists a torsion-free hyperbolic quotient $G\to G_{m}$, injective on $B$, of cohomological dimension $16$, such that for every unitary $G_{m}$-representation $V$, the second homology is $H_{2}(G_{m};V)\cong V^{m}$.
@@
The group $G$ is torsion-free hyperbolic of cohomological dimension $16$, and by Theorem 3.10 it has property $T_{2}]$. We construct $G_{m}$ by induction on $n$, starting with $G_{0}=G$, which works by Proposition 3.8.
Suppose by induction that $G_{m}$ has been constructed. Apply Proposition 4.13 with the finite set $B$ (and any $x,H$), let $g$ be the resulting element and $G_{m+1}$ the Dehn filling quotient. Hence $G_{m+1}$ is non-elementary hyperbolic, the quotient $G\to G_{m+1}$ is injective on $B$, and $(G_{m},\langle g\rangle,\langle g\rangle)$ is a Cohen–Lyndon triple. Now $G$ is the fundamental group of a closed $16$-dimensional aspherical manifold, so it has cohomological dimension $16$, and $H^{16}(G;\mathbb{Z})\cong\mathbb{Z}$. The first part of Corollary 4.3 therefore implies that the cohomological dimension of $G_{m+1}$ is $16$, hence in particular that $G_{m+1}$ is torsion-free. Moreover, since $G_{m}$ has property $(T)$, applying Proposition 3.8 and the second part of Corollary 4.3, we obtain, for every unitary $G_{m+1}$-representation $V$, a short exact sequence
$0\to H_{2}(G_{m};V)\to H_{2}(G_{m+1};V)\to V\to 0.$
By the induction hypothesis the first term is $V^{m}$. Therefore $H_{2}(G_{m+1};V)\cong V^{m+1}$, which proves the claim and hence the theorem. ∎
@@
The groups $G_{m}$ of course fail to have property $(T_{2})$, despite being quotients of the $(T_{2})$ group $G$, as remarked in [FF25]. Still, the second homology (and similarly the second cohomology) functor enjoys some strong form of rigidity: in a way, its output can be read directly from its input, and the same will be true in the next constructions. It would be interesting to explore whether a (co)homological rigidity of this form characterises quotients of $(T_{2})$ groups.
@@
We have $b^{(2)}_{2}(G)=0$ by Corollary 3.9. For each $i\in\mathbb{N}$, we let $B_{i}$ be the ball of radius $i$ in $G$, for a fixed choice of finite generating set, and $m_{i}\in\mathbb{N}$ as in the statement of C. Then Theorem 5.1 builds a group $G_{i}$, which by Proposition 3.6 has
$b^{(2)}_{2}(G_{i})=\dim_{\mathcal{N}(G_{i})}H_{2}(G_{i};\ell^{2}(G_{i}))=\dim_{\mathcal{N}(G_{i})}\ell^{2}(G_{i})^{m_{i}}=m_{i}.$
@@
In Theorem 5.1, we used the full power of property $T_{2}]$, but the proof shows that if $V$ is assumed to have no non-zero invariant vectors, then property $(T_{2})$ suffices. In the proof of C, we applied Theorem 5.1 only with $V=\ell^{2}(G_{i})$, which has no non-zero invariant vectors, so property $(T_{2})$ would have been sufficient. Similarly, the computation of second $\ell^{2}$-Betti numbers in the rest of this section and in the next could be obtained only using property $(T_{2})$.
5.2 Realising rational numbers
Next we consider arbitrary rational numbers, completing the proof of B.
Theorem 5.4 (corresponding to B).
Let $x\in\mathbb{Q}_{\geqslant 0}$ and let $\pi$ be the set of prime factors of the denominator of $x$. Then there exists a hyperbolic group $G$ with property $(T)$ and cohomological dimension $16$ modulo $\pi$, such that $b^{(2)}_{2}(G)=x$.
Proof.
Let $G_{0}$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, which has Euler characteristic $\chi\geqslant 1$ [Lüc02, Corollary 5.16(1) on p. 231]. If $x=0$, then we are done. Otherwise let $x=\frac{m}{k}$ be a reduced fraction, where $m,k\in\mathbb{N}$. We will build by induction a sequence of epimorphisms of hyperbolic groups $G_{0}\to G_{1}\to\cdots\to G_{m}$ and elements $g_{i}\in G_{i}$, such that:
  1. 1.
    $G_{i}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups, and cohomological dimension $16$ modulo $\pi$.
  2. 2.
    The element $g_{i}\in G_{i}$ has infinite order, and for all $j>i$ its image in $G_{j}$ has order $k$.
  3. 3.
    Each quotient $G_{i}\to G_{i+1}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{i+1}]$-modules and degree at least $3$, where $R$ is any ring where every prime from $\pi$ is invertible.
  4. 4.
    For every unitary $G_{i}$-representation $V$, the second homology is
    $H_{2}(G_{i};V)=\bigoplus_{j<i}V_{g_{j}}.$
The group $G=G_{m}$ is then non-elementary hyperbolic, with no non-trivial finite normal subgroups and cohomological dimension $16$ modulo $\pi$. It is a quotient of $G_{0}$ hence it has property $(T)$. Each $g_{i}:i<m$ has order $k$ in $G$, so by Propositions 3.6 and 3.11:
$\dim_{\mathcal{N}(G)}\ell^{2}(G)_{g_{i}}$ $=\dim_{\mathcal{N}(G)}H_{0}(\langle g_{i}\rangle,\ell^{2}(G))=\dim_{\mathcal{N}(G)}H_{0}(\langle g_{i}\rangle,\mathcal{N}(G))$
$=\dim_{\mathcal{N}(\langle g_{i}\rangle)}H_{0}(\langle g_{i}\rangle,\mathcal{N}(\langle g_{i}\rangle))=b^{(2)}_{0}(\langle g_{i}\rangle)=\frac{1}{k}.$
By Proposition 3.6 and additivity we therefore have
$b^{(2)}_{2}(G)=\dim_{\mathcal{N}(G)}H_{2}(G;\ell^{2}(G))=\sum\limits_{0\leqslant i<m}\frac{1}{k}=x.$
By Item 3, for all $n\geqslant 3$ we have
$b^{(2)}_{n}(G_{m})=\dim_{\mathcal{N}(G_{x})}H_{n}(G_{m};\mathcal{N}(G_{m}))=\dim_{\mathcal{N}(G_{m})}H_{n}(G_{0};\mathcal{N}(G_{m}))=b^{(2)}_{n}(G_{0}\to G_{x}).$
Since $G_{m}$ has property $(T)$, $b^{(2)}_{1}(G_{m})=0$ by [BV97, Corollary 6]. Hence
$\chi^{(2)}(G_{m})$ $=\sum\limits_{n\in\mathbb{N}}(-1)^{n}b^{(2)}_{n}(G_{m})=b^{(2)}_{2}(G_{m})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{m})$
$=b^{(2)}_{2}(G_{m})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{0}\to G_{m})=b^{(2)}_{2}(G_{m})+\chi,$
where the last equality follows from Corollary 3.9.
It remains to construct the sequence. Suppose by induction that $G_{i}$ has been constructed, for some $0\leqslant i<n$. Apply Proposition 4.13 with $k$ (and any $x,H$), let $g_{i}$ be the resulting element and $G_{i+1}$ the Dehn filling quotient; we also choose $B$ to be sufficiently large such that for all $j<i$, the element $g_{j}$ still has order $k$ in $G_{i+1}$. Hence $G_{i+1}$ is non-elementary hyperbolic with no non-trivial finite normal subgroup, and $(G_{i},\langle g_{i}\rangle,\langle g_{i}^{k}\rangle)$ is a Cohen–Lyndon triple. Induction and the first part of Corollary 4.3 imply that $G_{i+1}$ has cohomological dimension $16$ modulo $\pi$. Moreover, the second part and property $(T)$ give, for every unitary $G_{i+1}$-representation $V$, a short exact sequence
$0\to H_{2}(G_{i};V)\to H_{2}(G_{i+1};V)\to V_{g_{i}}\to 0.$
This concludes the construction, hence the proof. ∎
6 Reals and simple groups
@@
Let $p$ be a prime. There exists a family of groups $(G_{x})_{x\in\mathbb{R}_{>0}}$ and an integer $\chi\geq 1$ with the following properties.
@@
    $G_{x}$ is simple, has property $(T)$, and is lacunary hyperbolic.
@@
    $G_{x}$ has cohomological dimension $16$ modulo $p$.
@@
    The second $\ell^{2}$-Betti number is $b^{(2)}_{2}(G_{x})=x$.
@@
    The $\ell^{2}$-Euler characteristic is $\chi^{(2)}(G_{x})=b^{(2)}_{2}(G_{x})+\chi$.
@@
We fix a positive real $x\in\mathbb{R}_{>0}$ and choose a base-$p$ representation
$\sum\limits_{i\geqslant 1}\frac{1}{p^{k_{i}}},$
for some $k_{i}\in\mathbb{N}$. To treat all numbers in a unified way, if $x$ is $p$-adic, we still choose a periodic representation. The group $G_{x}$ will arise as the colimit of a direct sequence of epimorphisms of hyperbolic groups $G_{0}\to G_{1}\to\cdots$, where $G_{0}$ is a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane, which has Euler characteristic $\chi\geqslant 1$ [Lüc02, Corollary 5.16(1) on p. 231]. We fix a finite generating set $S_{0}$ for $G_{0}$ and let $S_{i}$ be the pushforward to $G_{i}$. We also specify elements $g_{i}\in G_{i}$, and will prove the following additional properties.
@@
    The element $g_{i}\in G_{i}$ has infinite order, and for all $j>i$ its image in $G_{j}$ has order $p^{k_{i}}$.
@@
Let us first prove that the colimit $G_{x}$ has the desired properties. The group $G_{0}$ has property $(T)$ by Theorem 3.10, hence so does $G_{x}$, being its quotient. Item 2 implies that $G_{x}$ is lacunary hyperbolic, by Theorem 4.14. Item 3 and Corollary 2.4 imply that $G_{x}$ has cohomological dimension $16$ modulo $p$.
Next, we compute $b^{(2)}_{2}(G_{x})$. By Proposition 2.2, we have
$H_{2}(G_{x};\ell^{2}(G_{x}))\cong\bigoplus_{i\geqslant 1}\ell^{2}(G_{x})_{g_{i}}.$
Each $g_{i}$ has order $p^{k_{i}}$ in $G_{x}$, hence by Propositions 3.6 and 3.11:
$\dim_{\mathcal{N}(G_{x})}\ell^{2}(G_{x})_{g_{i}}$ $=\dim_{\mathcal{N}(G_{x})}H_{0}(\langle g_{i}\rangle,\ell^{2}(G_{x}))=\dim_{\mathcal{N}(G_{x})}H_{0}(\langle g_{i}\rangle,\mathcal{N}(G_{x}))$
$=\dim_{\mathcal{N}(\langle g_{i}\rangle)}H_{0}(\langle g_{i}\rangle,\mathcal{N}(\langle g_{i}\rangle))=b^{(2)}_{0}(\langle g_{i}\rangle)=p^{-k_{i}}.$
@@
$b^{(2)}_{2}(G_{x})=\dim_{\mathcal{N}(G_{x})}H_{2}(G_{x};\ell^{2}(G_{x}))=\sum\limits_{i\geqslant 1}\frac{1}{p^{k_{i}}}=x.$
Finally, we compute $\chi^{(2)}(G_{x})$. By Item 3 and Proposition 2.2, for all $n\geqslant 3$ we have
$b^{(2)}_{n}(G_{x})=\dim_{\mathcal{N}(G_{x})}H_{n}(G_{x};\mathcal{N}(G_{x}))=\dim_{\mathcal{N}(G_{x})}H_{n}(G_{0};\mathcal{N}(G_{x}))=b^{(2)}_{n}(G_{0}\to G_{x}).$
Since $G_{x}$ has property $(T)$, $b^{(2)}_{1}(G_{x})=0$ by [BV97, Corollary 6]. Hence
$\chi^{(2)}(G_{x})$ $=\sum\limits_{n\in\mathbb{N}}(-1)^{n}b^{(2)}_{n}(G_{x})=b^{(2)}_{2}(G_{x})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{x})$
$=b^{(2)}_{2}(G_{x})+\sum\limits_{n\geqslant 3}(-1)^{n}b^{(2)}_{n}(G_{0}\to G_{x})=b^{(2)}_{2}(G_{x})+\chi,$
@@
It remains to construct the directed sequence $G_{0}\to G_{1}\to\cdots$ and the elements $g_{i}\in G_{i}$ so that the items above are satisfied. Starting with $G_{0}$, note that Item 5 holds by Proposition 3.8, because $G_{0}$ has property $T_{2}]$, and there is nothing to check for the other items, before $g_{0}$ has been chosen. So suppose by induction that $G_{0},\ldots,G_{i}$ and the elements $g_{0},\ldots,g_{i-1}$ have been constructed so that all items above are satisfied. Apply Proposition 4.13 with $k=k_{i}$, a finite set $B$ to be specified later (and any $x,H$). Let $g_{i}$ be the resulting element and $G_{i+1}$ the Dehn filling quotient. Proposition 4.13 directly gives Item 1. By choosing $B$ large enough, we can ensure that Item 2 is satisfied, and that the elements $g_{j}:j<i$ maintain the same order in $G_{i+1}$; this together with Proposition 4.13 gives Item 4.
Because $(G_{i},\langle g_{i}\rangle,\langle g_{i}^{k}\rangle)$ is Cohen–Lyndon, again by Proposition 4.13, we can apply Corollary 4.3, the first part gives Item 3. The second part, together with the fact that $G_{i}$ has property $(T)$, gives a short exact sequence
@@
We now prove A. The proof will be similar, but more technical, in that we need to leverage an additional prime, and the full power of Proposition 4.13, to make the group $G_{x}$ simple. This improvement employs a trick from [OT13], where it was used to construct simple groups with large first $\ell^{2}$-Betti numbers. The following will be our starting point, this is also the only place where we use Proposition 4.12 for the induction step.
@@
Let $G$ be a torsion-free cocompact lattice in the isometry group of the octonionic hyperbolic plane. Let $q$ be a prime and let $\varepsilon>0$. Then there exists a quotient $G\to G_{0}$ with the following properties.
@@
    $G_{0}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
@@
    The quotient $G\to G_{0}$ induces isomorphisms in homology and cohomology with coefficients in $R[G_{0}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible.
@@
    $G_{0}$ is generated by elements $\{a_{1},\ldots,a_{m}\}$, each of which has order a power of $q$.
@@
    There is a finite set $\mathcal{E}$ of finite subgroups of $G_{0}$, such that for every unitary $G_{0}$-representation $V$, the second homology is
    $H_{2}(G_{0};V)=\bigoplus_{E\in\mathcal{E}}V_{E},$
    and moreover
    $\sum\limits_{E\in\mathcal{E}}|E|^{-1}<\varepsilon.$
@@
We start by choosing a generating set $\{a_{1}^{0},\ldots,a_{m}^{0}\}$ of $G$. For $i=1,\ldots,m$, we construct by induction a quotient $G\to G^{i}$ with the following properties.
@@
    $G^{i}$ is non-elementary hyperbolic with no non-trivial finite normal subgroups.
@@
    The quotient $G\to G^{i}$ induces isomorphisms in homology and cohomology with coefficients in $R[G^{i}]$-modules and degree at least $3$, where $R$ is any ring where $q$ is invertible. In particular $G^{i}$ has cohomological dimension $16$ modulo $q$, and so it only contains $q$-torsion [Bie81, Proposition 4.11 on p. 63].
@@
    $G^{i}$ is generated by elements $\{a_{1}^{i},\ldots,a_{m}^{i}\}$, where $a_{1}^{i},\ldots,a_{i}^{i}$ have order a power of $q$.